OpenStudy (anonymous):

Five candidates are running for three different class offices: president, vice president, and secretary. In how many ways can the offices be filled with these candidates?

OpenStudy (neer2890):

\[\left(\begin{matrix}5 \\ 3\end{matrix}\right)\]

OpenStudy (kropot72):

The order of selection is important, therefore the number of permutations of the five candidates taken three at a time is needed. \[5P3=\frac{5!}{(5-3)!}=\frac{5\times4\times3\times2}{2}=you\ can\ calculate\]

OpenStudy (goformit100):

Warm Welcome To Open Study ......... Sir/ Ma'am @samanthataylee